OpenStudy (anonymous):

find the slope (-1,3) (4,-3)

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(-3-3)/(4-(-1)) m=-6/5

OpenStudy (anonymous):

y2-y1 over x2-x1 so -3-3 over 4-(-1) = -6/5

OpenStudy (anonymous):

sweet thanks guys :d